On a 5 hour trip, Thomas drove part of the time at 30 mi/h and the rest at 55 mi/h. If he traveled a distance of 250 mi, how much time did he spend driving at 55 mi/h?
Answer by mananth(16946)

On a 5 hour trip, Thomas drove part of the time at 30 mi/h and the rest at 55 mi/h. If he traveled a distance of 250 mi, how much time did he spend driving at 55 mi/h?
...
let the first leg distance traveled be xmiles
speed = 30mph
time = x/30 hours
...
next leg = 250-x miles
speed = 55 mph
time = 250-x / 55
...
x/30 + (250-x)/55 =5
30(250-x)+55x=5*30*55
7500-30x+55x=5*30*55
25x= 8250-7500
25x=750
x=30 miles
speed = 30mph
distance = 30 miles
time = 1 hour
so he drove 4 hours at 55mph
...
m.ananth@hotmail.ca